OpenStudy (anonymous):

A rectangular playground has a perimeter of 60m and an area of 216m^2. Find the dimensions of the playground. Thanks :)

OpenStudy (anonymous):

2(l+b)=60 l*b=216

OpenStudy (anonymous):

2l+2b=60 2l=60-2b l=30-b

OpenStudy (anonymous):

he's on the way yes?

OpenStudy (anonymous):

\[(30-b)b=216\] \[30b-b^2=216\] \[b^2-30b+216=0\] \[(b-18)(b-12)=0\] \[b=18\] or \[b=12\]

OpenStudy (anonymous):

so i am willing to be one side is 12 and the other is 18

OpenStudy (anonymous):

check is easy enough, \[18\times 12=216\] and \[18+12+18+12=60\]

OpenStudy (lalaly):

(2a+2b)=60 ab=216 a=216/b 2(216/b)+2b=60 b=18 or 12 substitute to find a a*18=216 a=12 a*12=216 a=18 the sides are 18 and 12

OpenStudy (anonymous):

thank u very much :) rly appreciated ur help!

OpenStudy (anonymous):

satellite, isn't 30b-b^2=216 equal to.. b^2 - 30b - 216 = 0 and not b^2 - 30b + 216 = 0 why must we add 216?

do any of u know why? ):